1. The profit of a company is given by P(x) = 5 000 +1 000x-5x² where x is the amount (in thousands of pesos) that the company spends on advertising.




a. Find the amount x that the company has to spend to maximize its profit.





b. Find the maximum profit.


1
Expert's answer
2021-12-14T08:16:53-0500
"P(x) = 5 000 +1 000x-5x^2, x\geq 0"

Find the first derivetive with respect to "x"


"P'(x)=1000-10x"

Find the critical number(s)


"P'(x)=0=>1000-10x=0"

"x=100"

If "0<x<100, P'(x)>0, P(x)" increases.

If "x>100, P'(x)<0, P(x)" decreases.

The function "P(x)" has a local maximum at "x=100."

Since the function "P" has the only extremum, then the function "P(x)" has the absolute maximum at "x=100."

a. The company has to spend "100,000" pesos has to maximize its profit.


b.


"P(100) = 5 000 +1 000(100)-5(100)^2"

"=55000"

The maximum profit is "55000."